# Faster Linear Systems and Matrix Norm Approximation via Multi-level Sketched Preconditioning

### v1

@@ 1.3 Additional Related Work
Also related to our work is remarkable recent progress on analyzing and improving stochastic iterative methods like stochastic gradient descent, stochastic coordinate descent, and variants thereof [SV09, RSB12, JZ13, SSZ14]. A major development in the area is that it is possible to obtain runtimes for solving linear systems that depend on the average condition number, $\bar{\kappa}=\frac{1}{n}\sum_{i=1}^{n}\frac{\sigma_{i}}{\sigma_{n}}$ instead of the standard condition number $\kappa=\frac{\sigma_{1}}{\sigma_{n}}$. E.g., for PSD linear systems, [LS13] presents a variant of coordinate descent that runs in $\tilde{O}(n^{2}\sqrt{\bar{\kappa}})$ time vs. $\tilde{O}(n^{2}\sqrt{{\kappa}})$ for the conjugate gradient method. Similar results have been proven for variants of stochastic gradient descent [FGKS15]. Naturally, the average condition $\bar{\kappa}$ can be significantly smaller than $\kappa$ if $\mathbf{A}$ only has a few large singular values and is otherwise well-conditioned. Perhaps unsurprisingly then, the current state-of-the-art methods for 1.1 and downstream applications like matrix norm approximation rely on stochastic iterative methods [MNS+18, DY24]. A high-level observation of our work is that it is possible to match and actually exceed the efficiency of these methods using deterministic iterative methods with randomized precondition. This will become more apparent in Section 3, where our main theorem is stated in terms of a dependence on an average condition number.
@@ 2 Preliminaries
Our results are proven in the real RAM model. We assume the inputs $\mathbf{A}$ and $\mathbf{b}$ are given with real-valued entries, and basic arithmetic operations ($+,-,\times,\div,\sqrt{\cdot}$) can be performed exactly on real numbers in $O(1)$ time. We believe that all methods described in this paper are numerically stable, i.e. that our results could be extended to a weaker finite precision model of arithmetic. However, doing so is beyond the scope of this work. We emphasize that we do leverage results on the numerical stability of iterative solvers like the Lanczos method in finite precision [Pai76, MMS18]. These results are used to account for the fact that we apply preconditioners inexactly, even in the real RAM model, not to address finite precision considerations.
@@ 3 Main Technical Results
Given $\mathbf{A}\in\mathbb{R}^{m\times n}$ with condition number $\kappa$, vector $\mathbf{b}\in\mathbb{R}^{n}$ and regularization parameter $\lambda\geq 0$, let $\{\sigma_{i}\}_{i=1}^{n}$ be the singular values of $\mathbf{A}$ in decreasing order, and let $\mathbf{x}^{*}=(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I})^{-1}\mathbf{c}$. For any $l\in(\log n+1,\ldots,n)$, define $\bar{\kappa}_{l,\lambda}\coloneqq(\frac{1}{n-l}\sum_{i>l}\sigma_{i}^{2}/(\sigma_{n}^{2}+\lambda))^{1/2}$. Given $\epsilon>0$ and $\delta\in(0,1/8)$, if Algorithm 3 is run with $\lambda_{0}=\frac{2}{l}\sum_{i>l}\sigma_{i}^{2},s=O(l\log(l/\delta)),\gamma=O(\log(l/\delta))$ and $t_{\max}=O(\sqrt{n/l}\cdot\bar{\kappa}_{l,\lambda}\log(\bar{\kappa}_{l,\lambda}/\epsilon))$, then with probability at least $1-\delta$, it will output $\tilde{\mathbf{x}}$ such that $\|\tilde{\mathbf{x}}-\mathbf{x}^{*}\|_{\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I}}\leq\epsilon\|\mathbf{x}^{*}\|_{\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I}}$ in time
@@
We next demonstrate how Theorem 3.1 can be used to show our main result for matrices with $k$ large singular values. To do so, we will optimize over the choice of $l$
@@
It remains to obtain the desired time complexity by optimizing over $l$, which we will choose to be larger than $k$. Notice that for $l>k$, we have $\bar{\kappa}_{l,0}\leq\sigma_{l+1}/\sigma_{n}=O(1)$ under the assumption of Theorem 1.1. So, upper bounding $\mathrm{nnz}(\mathbf{A})\leq n^{2}$, we can upper bound the runtime in Theorem 3.1 by:
$\tilde{O}(n^{2.5}/\sqrt{l}+l^{\omega}),\quad\text{for any $l\geq k$}.$
@@
  • If $k\leq n^{5/(2\omega+1)}$, then we choose $l=n^{5/(2\omega+1)}$ and obtain time complexity $\tilde{O}(l^{\omega})=\tilde{O}(n^{2+\frac{\omega-2}{2\omega+1}})$.
  • If $n^{5/(2\omega+1)}<k<n$, then we choose $l=k$, and obtain time complexity $\tilde{O}(l^{\omega})=\tilde{O}(k^{\omega})$.
@@
$\tilde{O}(n^{2+\frac{\omega-2}{2\omega+1}}+k^{\omega}),$
which, using current matrix multiplication exponent $\omega\approx 2.372$, simplifies to $\tilde{O}(n^{2.065}+k^{\omega})$. ∎
@@ 3.1 Applications to Regularized Linear Systems and Least Squares
This result is obtained by combining our MSP method with a standard sketch-and-precondition iterative solver [RT08, Epp24]. In particular, for a tall matrix $\mathbf{A}\in\mathbb{R}^{n\times d}$ where $n\gg d$, one can first construct a sketch $\tilde{\mathbf{A}}$ with $O(d)$ rows using a constant-factor sparse subspace embedding (e.g., Lemma 4.4) which takes $\tilde{O}(\mathrm{nnz}(\mathbf{A})+d^{2})$ time. To solve a least squares problem involving $\mathbf{A}$, it suffices to solve $O(\log(1/\epsilon))$ linear systems of the form $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})^{-1}\mathbf{g}$ for some vector $\mathbf{g}$ up to constant error. See e.g. Section 10 of [DY24] for details.
@@ 4 Two-Level MSP for Positive Definite Linear Systems
The construction of preconditioner $\mathbf{M}$ is based on Nyström approximation. Let $\mathbf{C}=\mathbf{A}\mathbf{S}^{\top}$ and $\mathbf{W}=\mathbf{S}\mathbf{A}\mathbf{S}^{\top}$ where $\mathbf{S}\in\mathbb{R}^{s\times n}$ is some sketching matrix, then the Nyström approximation of $\mathbf{A}$ can be expressed as $\hat{\mathbf{A}}_{\mathrm{nys}}=\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}$. We show that if we choose $\mathbf{S}$ to be the sparse embedding matrix with $s=\tilde{O}(l)$ rows and $\gamma=\tilde{O}(1)$ non-zero entries per column, then we have the approximation guarantee $\|\hat{\mathbf{A}}_{\mathrm{nys}}-\mathbf{A}\|\leq\frac{2}{l}\sum_{i>l}\lambda_{i}$, which in turn gives $\kappa_{\mathbf{M}}=O(\bar{\kappa}_{l,\lambda}n/l)$ for $\mathbf{M}\coloneqq\hat{\mathbf{A}}_{\mathrm{nys}}+\tilde{\lambda}\mathbf{I}$ where $\tilde{\lambda}\coloneqq\lambda+\frac{2}{l}\sum_{i>l}\lambda_{i}$, see Lemma 4.2. With the choice of $\mathbf{M}$ we can verify the following holds:
@@
Let $\mathbf{\Phi}\in\mathbb{R}^{\phi\times n}$ be the sparse embedding matrix in Lemma 4.4 with $\phi=\tilde{O}(s)$, with the guarantee of subspace embedding, we can construct the Level-2 preconditioner as $\mathbf{M}_{2}\coloneqq\tilde{\mathbf{C}}^{\top}\tilde{\mathbf{C}}+\lambda\mathbf{W}$ where $\tilde{\mathbf{C}}=\mathbf{\Phi}\mathbf{C}$ in time $\tilde{O}(ns+s^{3})$. Notice that this step only needs to be done once. By transforming the error bound (that we get from Level-2 preconditioned Lanczos) from $\hat{\mathbf{y}}$ to $\hat{\mathbf{w}}$ using Lemma 4.3, we can verify the assumption Eq.(4.1) which is required in Level-1 preconditioned Lanczos, thus finishing the proof. Formally, we have the following theorem for solving positive definite linear systems, where note that the condition number $\bar{\kappa}_{l,\lambda}$ is defined differently in terms of the eigenvalues of $\mathbf{A}$, rather than in terms of the squared singular values. Also note that although we assume $\mathbf{A}\in\mathcal{S}_{n}^{++}$, indeed we only need $\mathbf{A}\in\mathcal{S}_{n}^{+}$ and $\mathbf{A}+\lambda\mathbf{I}\in\mathcal{S}_{n}^{++}$ (which is satisfied if $\lambda>0$).
@@
Given $\mathbf{A}\in\mathcal{S}_{n}^{++}$ with condition number $\kappa$, $\mathbf{b}\in\mathbb{R}^{n}$ and regularized term $\lambda\geq 0$. Let $\{\lambda_{i}\}_{i=1}^{n}$ be the eigenvalues of $\mathbf{A}$ in decreasing order, let $\mathbf{x}^{*}=(\mathbf{A}+\lambda\mathbf{I})^{-1}\mathbf{b}$ be the solution of the regularized linear system $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$. For any $\log n<l<n$, we define $\bar{\kappa}_{l,\lambda}\coloneqq\frac{1}{n-l}\sum_{i>l}^{n}\lambda_{i}/(\lambda_{n}+\lambda)$. Given $\epsilon>0$ and $80<\delta<1/$ 8, with probability at least $1-\delta$, running Algorithm 1 with choice $\lambda_{0}=\frac{2}{l}\cdot\sum_{i>l}\lambda_{i},s=O(l\log(l/\delta)),\gamma=O(\log(l/\delta))$ and $t_{\max}=O(\sqrt{\bar{\kappa}_{l,\lambda}n/l}\cdot\log(\bar{\kappa}_{l,\lambda}n/\epsilon))$ will output $\tilde{\mathbf{x}}$ such that $\|\tilde{\mathbf{x}}-\mathbf{x}^{*}\|_{\mathbf{A}+\lambda\mathbf{I}}\leq\epsilon\|\mathbf{x}^{*}\|_{\mathbf{A}+\lambda\mathbf{I}}$ in time
@@
1:Input: matrix $\mathbf{A}\in\mathcal{S}_{n}^{++}$, regularize term $\lambda$ and $\lambda_{0}$, vector $\mathbf{b}\in\mathbb{R}^{n}$, sparse sketch size $s$, # of non-zeros $\gamma$, # of iterations $t_{\max}$;
2:Compute $\tilde{\lambda}=\lambda+\lambda_{0}$, $\mathbf{C}=\mathbf{A}\mathbf{S}^{\top}$, $\mathbf{W}=\mathbf{S}\mathbf{C}$;
3:Compute $\tilde{\mathbf{C}}=\mathbf{\Phi}\mathbf{C}$ and $\mathbf{M}_{2}^{-1}=(\tilde{\mathbf{C}}^{\top}\tilde{\mathbf{C}}+\tilde{\lambda}\mathbf{W})^{-1}$; $\triangleright$ Lemma 4.4.
4:Solve $\tilde{\mathbf{x}}$ by calling preconditioned Lanczos with $(\mathbf{A}+\lambda\mathbf{I},\mathbf{b},\mathsf{SolveM1},t_{\max})$; $\triangleright$ Alg. 6.
5:return $\tilde{\mathbf{x}}$; $\triangleright$ Solves $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$.
@@
Let $\mathbf{S}\in\mathbb{R}^{s\times n}$ be the sparse embedding matrix according to Lemma 4.2 with $s=O(l\log(l/\delta))$ and $\gamma=O(\log(l/\delta))$. Denote $\mathbf{C}\coloneqq\mathbf{A}\mathbf{S}^{\top}$ and $\mathbf{W}\coloneqq\mathbf{S}\mathbf{A}\mathbf{S}^{\top}$, then we have $\hat{\mathbf{A}}_{\mathrm{nys}}=\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}$ and $\mathbf{M}=\hat{\mathbf{A}}_{\mathrm{nys}}+\tilde{\lambda}\mathbf{I}$ where $\tilde{\lambda}=\lambda+\lambda_{0}$. According to Lemma 4.2, with probability at least $21-\delta/$ 2 we have $\kappa_{\mathbf{M}}\leq\frac{C\bar{\kappa}_{l,\lambda}n}{l}$ for some $C=O(1)$.
@@
Given positive definite matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ and $\delta\in(0,1/2)$. For $\log n<l<n$, let $\hat{\mathbf{A}}_{\mathrm{nys}}=\mathbf{A}\mathbf{S}^{\top}(\mathbf{S}\mathbf{A}\mathbf{S}^{\top})^{-1}\mathbf{S}\mathbf{A}$ be the Nyström approximation of $\mathbf{A}$, where $\mathbf{S}\in\mathbb{R}^{s\times n}$ is the sparse embedding matrix with $s=O(l\log(l/\delta))$ and $\gamma=O(\log(l/\delta))$. Given $\lambda\geq 0$, let $\mathbf{M}=\hat{\mathbf{A}}_{\mathrm{nys}}+(\lambda+\lambda_{0})\mathbf{I}$ where $\lambda_{0}=\frac{2}{l}\sum_{i\geq l+1}\lambda_{i}$, then with probability $1-\delta$ we have
@@
For the first level we use preconditioned Lanczos (Algorithm 6) to solve $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$ with preconditioner $\mathbf{M}$, where we need to compute $\mathbf{M}^{-1}\mathbf{r}$ in each iteration. The following lemma guarantees that by using inverse formula, we can approximate $\mathbf{M}^{-1}\mathbf{r}$ well enough by solving a linear system with matrix $\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W}$ to a certain accuracy.
@@
Denote $\hat{\mathbf{w}}=\frac{1}{\tilde{\lambda}}(\mathbf{r}-\mathbf{C}\hat{\mathbf{y}})$, then $\hat{\mathbf{w}}$ is an approximate solution of the linear system $\mathbf{M}\mathbf{w}=\mathbf{r}$ satisfying
@@
By using this inverse formula, we have $\mathbf{w}^{*}\coloneqq\mathbf{M}^{-1}\mathbf{r}=\frac{1}{\tilde{\lambda}}(\mathbf{r}-\mathbf{C}\mathbf{y}^{*})$. Instead of directly computing $\mathbf{y}^{*}$, suppose we can solve the linear system $(\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W})\mathbf{y}=\mathbf{C}^{\top}\mathbf{r}$ approximately and have the guarantee (which will be proved in Level-2 analysis later) that $\|\hat{\mathbf{y}}-\mathbf{y}^{*}\|_{\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W}}\leq\epsilon_{1}\cdot\|\mathbf{y}^{*}\|_{\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W}}$, then Lemma 4.3 guarantees that Eq.(4.2) hold as long as $\mathbf{M}\approx_{2}\mathbf{A}+\tilde{\lambda}\mathbf{I}$. The assumption here can be easily verified: for one side we have

### v2

@@ 1.3 Additional Related Work
Also related to our work is recent progress on analyzing and improving stochastic iterative methods like stochastic gradient descent, randomized Kaczmarz, stochastic coordinate descent, and variants thereof [SV09, RSB12, JZ13, SSZ14]. A major development in the area is that it is possible to obtain runtimes for solving linear systems that depend on the average condition number, $\bar{\kappa}=\frac{1}{n}\sum_{i=1}^{n}\frac{\sigma_{i}}{\sigma_{n}}$ instead of the standard condition number $\kappa=\frac{\sigma_{1}}{\sigma_{n}}$. E.g., for PD linear systems, [LS13] presents a variant of coordinate descent that runs in $\tilde{O}(n^{2}\sqrt{\bar{\kappa}})$ time vs. $\tilde{O}(n^{2}\sqrt{{\kappa}})$ for the conjugate gradient method. Similar results have been proven for variants of stochastic gradient descent [FGKS15]. More recently, [DLNR24] uses a connection between block coordinate descent and low-rank approximation [DR24], obtaining an algorithm that runs in $\tilde{O}(n^{2}\sqrt{\bar{\kappa}_{\ell}})$ time for $\ell=O(n^{\frac{1}{\omega-1}})$, where $\bar{\kappa}_{\ell}=\frac{1}{n-\ell}\sum_{i>\ell}\frac{\sigma_{i}}{\sigma_{n}}\leq\bar{\kappa}$. Naturally, these average condition numbers $\bar{\kappa}$ and $\bar{\kappa}_{\ell}$ can be significantly smaller than $\kappa$ if $\mathbf{A}$ only has a few large singular values and is otherwise well-conditioned. Perhaps unsurprisingly then, the current state-of-the-art methods for 1.1 and downstream applications like matrix norm approximation rely on stochastic iterative methods [MNS+18, DY24, DLNR24]. A high-level observation of our work is that it is possible to match and actually exceed the efficiency of these methods using deterministic iterative methods with randomized preconditioning. This will become more apparent in Section 3, where our main theorem is stated in terms of a dependence on an average condition number.
@@ 2 Preliminaries
Sparse Subspace Embedding.
Following [CNW16, Coh16], we define the sparse embedding matrix $\mathbf{S}\in\mathbb{R}^{s\times n}$ with sparsity parameter $\gamma$, which will be used in our construction of the Nyström preconditioner, as follows. Notice that given $\mathbf{A}\in\mathbb{R}^{n\times n}$, $\mathbf{S}\mathbf{A}$ can be computed in time $O(\gamma\cdot\mathrm{nnz}(\mathbf{A}))$.
Definition 2.1 (Sparse embedding matrix).
We define an $s\times n$ matrix $\mathbf{S}$ to be a sparse embedding matrix with sparsity $\gamma$, if the columns of $\mathbf{S}$ are independent, and for each column of $\mathbf{S}$, there are $\gamma$ random entries chosen uniformly without replacement and set to ±$\pm 1/\sqrt{\gamma}$ independently, with other entries in that column being set to 00.
Based on the construction of sparse embedding matrix, recent work [CDDR24] further constructs a fast oblivious subspace embedding (OSE) matrix and shows an optimal $O(d)$ result for embedding dimension, as stated in the following lemma. Throughout this paper, we will use $\mathbf{\Phi}$ to denote this oblivious embedding matrix, which will be used in the construction of a preconditioner for the second level of linear system solving.
Lemma 2.2 (Adapted from Theorem 1.4 in [CDDR24]).
@@
Our results are proven in the real RAM model (i.e, exact arithmetic). We assume the inputs $\mathbf{A}$ and $\mathbf{b}$ are given with real-valued entries, and basic arithmetic operations ($+,-,\times,\div,\sqrt{\cdot}$) can be performed exactly on real numbers in $O(1)$ time. We still need to leverage results on the finite-precision behavior of iterative solvers like the Lanczos method [Pai76, MMS18] to account for the fact that we apply preconditioners inexactly, even in the real RAM model. A full analysis of our methods in a finite precision model of computation is beyond the scope of this work, but is an important next step for future work on Multi-level Sketched Preconditioning.
@@ 3 Main Technical Results
Given $\mathbf{A}\in\mathbb{R}^{m\times n}$ with condition number $\kappa$, vector $\mathbf{b}\in\mathbb{R}^{n}$ and regularization parameter $\lambda\geq 0$, let $\{\sigma_{i}\}_{i=1}^{n}$ be the singular values of $\mathbf{A}$ in decreasing order, and let $\mathbf{x}^{*}=(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I})^{-1}\mathbf{c}$. For any $l\in\{\log n+1,\ldots,n\}$, define $\bar{\kappa}_{l,\lambda}\coloneqq(\frac{1}{n-l}\sum_{i>l}\sigma_{i}^{2}/(\sigma_{n}^{2}+\lambda))^{1/2}$. Given $\epsilon>0$ and $\delta\in(0,1/8)$, if Algorithm 3 is run with $\lambda_{0}=\frac{2}{l}\sum_{i>l}\sigma_{i}^{2},s=O(l\log(l/\delta)),\gamma=O(\log(l/\delta)),\phi=O(s+\log 1/\delta)$ and $t_{\max}=O(\sqrt{n/l}\cdot\bar{\kappa}_{l,\lambda}\log(\bar{\kappa}_{l,\lambda}/\epsilon))$, then with probability at least $1-\delta$, it will output $\tilde{\mathbf{x}}$ such that $\|\tilde{\mathbf{x}}-\mathbf{x}^{*}\|_{\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I}}\leq\epsilon\|\mathbf{x}^{*}\|_{\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I}}$ in time
@@
We next demonstrate how Theorem 3.1 can be used to show our main result for matrices with $k$ large singular values. To do so, we will optimize over the choice of $l$ as follows.
@@
It remains to obtain the desired time complexity by optimizing over $l$, which we will choose to be larger than $k$. Notice that for $l>k$, we have $\bar{\kappa}_{l,0}\leq\sigma_{l+1}/\sigma_{n}=O(1)$ under the assumption of Theorem 1.1. So, upper bounding $\mathrm{nnz}(\mathbf{A})\leq n^{2}$, we can upper bound the runtime in Theorem 3.1 by (here $\tilde{O}(\cdot)$ hide the logarithmic dependencies on $n$ and $\kappa$):
$\tilde{O}(n^{2.5}/\sqrt{l}\cdot\log^{3}1/\epsilon+l^{\omega}),\quad\text{for any $l\geq k$}.$
@@
  • If $k\leq n^{5/(2\omega+1)}$, we choose $l=n^{5/(2\omega+1)}$ and obtain time complexity $\tilde{O}(n^{2+\frac{\omega-2}{2\omega+1}}\log^{3}1/\epsilon)$.
  • If $n^{5/(2\omega+1)}<k<n$, we choose $l=k$, and obtain time complexity $\tilde{O}(n^{2+\frac{\omega-2}{2\omega+1}}\log^{3}1/\epsilon+k^{\omega})$.
@@
$\tilde{O}(n^{2+\frac{\omega-2}{2\omega+1}}\log^{3}1/\epsilon+k^{\omega}),$
which simplifies to $\tilde{O}(n^{2.065}\log^{3}1/\epsilon+k^{\omega})$ with current matrix multiplication exponent $\omega\approx 2.372$. ∎
@@ 3.1 Applications to Regularized Linear Systems and Least Squares
This result is obtained by combining our MSP method with a standard sketch-and-precondition iterative solver [RT08, Epp24]. In particular, for a tall matrix $\mathbf{A}\in\mathbb{R}^{n\times d}$ where $n\gg d$, one can first construct a sketch $\tilde{\mathbf{A}}$ with $O(d)$ rows using a constant-factor sparse oblivious subspace embedding (e.g., Lemma 2.2) which takes $\tilde{O}(\mathrm{nnz}(\mathbf{A})+d^{2})$ time. To solve a least squares problem involving $\mathbf{A}$, it suffices to solve $O(\log(1/\epsilon))$ linear systems of the form $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})^{-1}\mathbf{g}$ for some vector $\mathbf{g}$ up to constant error. See e.g. Section 10 of [DY24] for details.
@@ 4 Two-Level MSP for Positive Definite Linear Systems
The construction of preconditioner $\mathbf{M}$ is based on Nyström approximation. Let $\mathbf{C}=\mathbf{A}\mathbf{S}^{\top}$ and $\mathbf{W}=\mathbf{S}\mathbf{A}\mathbf{S}^{\top}$ where $\mathbf{S}\in\mathbb{R}^{s\times n}$ is some sketching matrix. Then the Nyström approximation of $\mathbf{A}$ can be expressed as $\hat{\mathbf{A}}_{\mathrm{nys}}=\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}$. We show that if we choose $\mathbf{S}$ to be a sparse embedding matrix (according to Definition 2.1) with $s=\tilde{O}(l)$ rows and $\gamma=\tilde{O}(1)$ non-zero entries per column, then we have the approximation guarantee $\|\hat{\mathbf{A}}_{\mathrm{nys}}-\mathbf{A}\|\leq\frac{2}{l}\sum_{i>l}\lambda_{i}$, which in turn gives $\kappa_{\mathbf{M}}=O(\bar{\kappa}_{l,\lambda}n/l)$ for $\mathbf{M}\coloneqq\hat{\mathbf{A}}_{\mathrm{nys}}+\tilde{\lambda}\mathbf{I}$ where $\tilde{\lambda}\coloneqq\lambda+\frac{2}{l}\sum_{i>l}\lambda_{i}$, see Lemma 4.2. With the choice of $\mathbf{M}$ we can verify the following holds:
@@
Let $\mathbf{\Phi}\in\mathbb{R}^{\phi\times n}$ be the oblivious subspace embedding matrix from Lemma 2.2 with $\phi=\tilde{O}(s)$. We can construct the Level-2 preconditioner as $\mathbf{M}_{2}\coloneqq\tilde{\mathbf{C}}^{\top}\tilde{\mathbf{C}}+\lambda\mathbf{W}$ where $\tilde{\mathbf{C}}=\mathbf{\Phi}\mathbf{C}$ in time $\tilde{O}(ns+s^{\omega})$. Notice that this step only needs to be done once. By transforming the error bound (that we get from Level-2 preconditioned Lanczos) from $\hat{\mathbf{y}}$ to $\hat{\mathbf{w}}$ using Lemma 4.3, we can verify the assumption Eq.(4.1) which is required in Level-1 preconditioned Lanczos, thus finishing the proof. Formally, we have the following theorem for solving positive definite linear systems, where note that the condition number $\bar{\kappa}_{l,\lambda}$ is defined differently in terms of the eigenvalues of $\mathbf{A}$, rather than in terms of the squared singular values. Also note that although we assume $\mathbf{A}\in\mathcal{S}_{n}^{++}$, indeed we only need $\mathbf{A}\in\mathcal{S}_{n}^{+}$ and $\mathbf{A}+\lambda\mathbf{I}\in\mathcal{S}_{n}^{++}$ (which is satisfied if $\lambda>0$).
@@
Given $\mathbf{A}\in\mathcal{S}_{n}^{++}$ with condition number $\kappa$, $\mathbf{b}\in\mathbb{R}^{n}$ and regularized term $\lambda\geq 0$. Let $\{\lambda_{i}\}_{i=1}^{n}$ be the eigenvalues of $\mathbf{A}$ in decreasing order, let $\mathbf{x}^{*}=(\mathbf{A}+\lambda\mathbf{I})^{-1}\mathbf{b}$ be the solution of the regularized linear system $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$. For any $\log n<l<n$, we define $\bar{\kappa}_{l,\lambda}\coloneqq\frac{1}{n-l}\sum_{i>l}^{n}\lambda_{i}/(\lambda_{n}+\lambda)$. Given $\epsilon>0$ and $\delta<1/8$, with probability at least $1-\delta$, running Algorithm 1 with choice $\lambda_{0}=\frac{2}{l}\cdot\sum_{i>l}\lambda_{i},s=O(l\log(l/\delta)),\gamma=O(\log(l/\delta)),\phi=O(s+\log 1/\delta)$ and $t_{\max}=O(\sqrt{\bar{\kappa}_{l,\lambda}n/l}\cdot\log(\bar{\kappa}_{l,\lambda}n/\epsilon))$ will output $\tilde{\mathbf{x}}$ such that $\|\tilde{\mathbf{x}}-\mathbf{x}^{*}\|_{\mathbf{A}+\lambda\mathbf{I}}\leq\epsilon\|\mathbf{x}^{*}\|_{\mathbf{A}+\lambda\mathbf{I}}$ in time
@@
1:Input: matrix $\mathbf{A}\in\mathcal{S}_{n}^{++}$, regularizer terms $\lambda$ and $\lambda_{0}$, vector $\mathbf{b}\in\mathbb{R}^{n}$, sparse sketch size $s$, # of non-zeros $\gamma$, level-2 sketch size $\phi$, # of iterations $t_{\max}$;
2:Construct sparse embedding matrix $\mathbf{S}\in\mathbb{R}^{s\times n}$ with $\gamma$ non-zeros per column; $\triangleright$ Def. 2.1.
3:Compute $\tilde{\lambda}=\lambda+\lambda_{0}$, $\mathbf{C}=\mathbf{A}\mathbf{S}^{\top}$, $\mathbf{W}=\mathbf{S}\mathbf{C}$;
4:Compute $\tilde{\mathbf{C}}=\mathbf{\Phi}\mathbf{C}\in\mathbb{R}^{\phi\times s}$ and $\mathbf{M}_{2}^{-1}=(\tilde{\mathbf{C}}^{\top}\tilde{\mathbf{C}}+\tilde{\lambda}\mathbf{W})^{-1}$; $\triangleright$ Lemma 2.2.
5:Solve $\tilde{\mathbf{x}}$ by calling preconditioned Lanczos with $(\mathbf{A}+\lambda\mathbf{I},\mathbf{b},\mathsf{SolveM1},t_{\max})$; $\triangleright$ Alg. 6.
6:return $\tilde{\mathbf{x}}$; $\triangleright$ Solves $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$.
@@
Let $\mathbf{S}\in\mathbb{R}^{s\times n}$ be the sparse embedding matrix according to Definition 2.1 with $s=O(l\log(l/\delta))$ and $\gamma=O(\log(l/\delta))$. Denote $\mathbf{C}\coloneqq\mathbf{A}\mathbf{S}^{\top}$ and $\mathbf{W}\coloneqq\mathbf{S}\mathbf{A}\mathbf{S}^{\top}$, then we have $\hat{\mathbf{A}}_{\mathrm{nys}}=\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}$ and $\mathbf{M}=\hat{\mathbf{A}}_{\mathrm{nys}}+\tilde{\lambda}\mathbf{I}$ where $\tilde{\lambda}=\lambda+\lambda_{0}$. According to Lemma 4.2, with probability at least $21-\delta/$ 2 we have $\kappa_{\mathbf{M}}\leq\frac{C\bar{\kappa}_{l,\lambda}n}{l}$ for some $C=O(1)$. For the proof of Lemma 4.2 and a detailed discussion of sparse embedding matrices, see Section 4.2.
@@
Given positive definite matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ and $\delta\in(0,1/2)$. For $\log n<l<n$, let $\hat{\mathbf{A}}_{\mathrm{nys}}=\mathbf{A}\mathbf{S}^{\top}(\mathbf{S}\mathbf{A}\mathbf{S}^{\top})^{-1}\mathbf{S}\mathbf{A}$ be the Nyström approximation of $\mathbf{A}$, where $\mathbf{S}\in\mathbb{R}^{s\times n}$ is the sparse embedding matrix with $s=O(l\log(l/\delta))$ and $\gamma=O(\log(l/\delta))$. Given $\lambda\geq 0$, let $\mathbf{M}=\hat{\mathbf{A}}_{\mathrm{nys}}+(\lambda+\lambda_{0})\mathbf{I}$ where $\lambda_{0}=\frac{2}{l}\sum_{i\geq l+1}\lambda_{i}$, then with probability $1-\delta$ the condition number $\kappa_{\mathbf{M}}$ of the matrix $\mathbf{M}^{-1/2}\mathbf{A}\mathbf{M}^{-1/2}$ satisfies:
@@
For the first level we use preconditioned Lanczos (Algorithm 6) to solve $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$ with preconditioner $\mathbf{M}$, where we need to compute $\mathbf{M}^{-1}\mathbf{r}$ in each iteration. The following lemma guarantees that by using inverse formula, we can approximate $\mathbf{M}^{-1}\mathbf{r}$ well enough by solving a linear system with matrix $\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W}$ to a certain accuracy. For the proof of Lemma 4.3, see Section 4.3.
@@
Then, $\hat{\mathbf{w}}=\frac{1}{\tilde{\lambda}}(\mathbf{r}-\mathbf{C}\hat{\mathbf{y}})$ is an approximate solution of the linear system $\mathbf{M}\mathbf{w}=\mathbf{r}$ satisfying
@@
By using this inverse formula, we have $\mathbf{w}^{*}\coloneqq\mathbf{M}^{-1}\mathbf{r}=\frac{1}{\tilde{\lambda}}(\mathbf{r}-\mathbf{C}\mathbf{y}^{*})$. Instead of directly computing $\mathbf{y}^{*}$, suppose we can solve the linear system $(\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W})\mathbf{y}=\mathbf{C}^{\top}\mathbf{r}$ approximately and have the guarantee (which will be proved in Level-2 analysis later) that $\|\hat{\mathbf{y}}-\mathbf{y}^{*}\|_{\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W}}\leq\epsilon_{1}\cdot\|\mathbf{y}^{*}\|_{\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W}}$, then Lemma 4.3 guarantees that Eq.(4.2) holds as long as $\mathbf{M}\approx_{2}\mathbf{A}+\tilde{\lambda}\mathbf{I}$. The assumption here can be easily verified: for one side we have